40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS
|
Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS   
|
Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS   
|
Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled
|
Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement  
|
Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement
|
Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened
|
Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement
|
Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement
|
Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement
|
|
Message ID: 13204
Entry time: Mon Aug 14 16:24:09 2017
In reply to: 13177
Reply to this: 13222
|
Author: |
gautam |
Type: |
Update |
Category: |
ALS |
Subject: |
Fiber ALS |
|
|
Today, I borrowed the fiber microscope from Johannes and took a look at the fibers coupled to the PDs. The PD labelled "BEAT PD AUX Y" has an end that seems scratched (Attachments #1 and #2). The scratch seems to be on (or at least very close to) the core. The other PD (Attachments #3 and #4) doesn't look very clean either, but at least the area near the core seems undamaged. The two attachments for each PD corresponds to the two available lighting settings on the fiber microscope.
I have not attempted to clean them yet, though I have also borrowed the cleaning supplies to facilitate this from Johannes. I also plan to inspect the ends of all other fiber connections before re-installing them.
Quote: |
Last week, we were talking about reviving the Fiber ALS box. Right now, it's not in great shape. Some changes to be made:
- Supply power to the PDs (Menlo FPD310) via a power regulator board. The datasheet says the current consumption per PD is 250 mA. So we need 500mA. We have the D1000217 power regulator board available in the lab. It uses the LM2941 and LM2991 power regulator ICs, both of which are rated for 1A output current, so this seems suitable for our purposes. Thoughts?
- Install power decoupling capacitors on the PDs.
- Clean up the fiber arrangement inside the box.
- Install better switches, plus LED indicators.
- Cover the box.
- Install it in a better way on the PSL table. Thoughts? e.g. can we mount the unit in some electronics rack and route the fibers to the rack? Perhaps the PSL IR and one of the arm fibers are long enough, but the other arm might be tricky.
Previous elog thread about work done on this box: elog11650
|
|
|
|
|
|
|
|